• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name C.Hyde-Wright

    Log entry time 00:20:31 on June11,2002

    Entry number 85821

    keyword=shift summary

    16:00 Run 1388 Kinematics Q1_L40
    one more hour of statistics will achieve 4%
    17:00 End 1390 End of Q1_L40
    decrease beam to 10 muA for Left Arm singles study
    17:10 CODA crash
    Kinematics Q1_L50
    HRS-R floor mark 82.5 vernier -4.5 = 82.671deg
    trig setup

    20:30 beam loss problems, current cut back to 80muA
    20:00 Gun problems EES called.